Quantitative Problem: Assume that interest rate parity holds. In the spot market 1 Japanese yen = $0.009, while in the 180-day forward market 1 Japanese yen = $0.0092. 180-day risk-free securities yield 1.05% in Japan. What is the yield on 180-day risk-free securities in the United States? Do not round intermediate calculations. Round your answer to two decimal places.

  %
Answer #1

Rate of Japan =1.05%
Spot market of Japanese Yen =0.009
Future market of Japanese Yen =0.0092
Future Rate =Spot Rate*(1+180 day US risk free Rate)5/(1+180 day Japan Risk free rate)
0.0092 =0.0090*(1+US Rate)/(1+1.05%)^0.5
(1+US Rate) =0.0092/0.0090*(1+1.05%) =1.03296
180 day US Rate =1.03295-1 =3.30%